# EgoTextVQA: Towards Egocentric Scene-Text Aware Video Question Answering

### v1

@@ 1 Introduction
The advances in wearable cameras and egocentric vision research [15, 26, 37], has led to a surge of interest in developing assistance applications. A promising direction is egocentric video question answering (VideoQA) [13, 52, 6, 12, 56], where the AI agents offer live help by answering user questions. Despite the potential, related dataset researches [54, 13, 20, 31, 10] have been designed to assess the VideoQA models in their visual understanding capabilities. The questions are not curated to reflect real-user needs in assistive applications (see EgoTaskQA in Fig. 1). QAEgo4D [2] and AssistQ [50] are closer in aim, but these datasets are designed to assist with episodic memory and “how to” demonstration; they emphasize pure visual understanding while neglecting scene text.
Research in assistive techniques [16, 42] have shown that scene text, including signs, labels, and other text elements that are ubiquitous in our daily lives, is frequently involved in user questions when seeking assistance. Existing efforts for QA on scene-text [42, 3] primarily aim at Optical Character Recognition (OCR) from ideal imaging conditions. They assume that people take good pictures and ask questions directly pinpointing the scene-text regions (see TextVQA [42] in Fig. 1). Despite the progress (e.g., leading results have accuracies of 85% on ST-VQA [3, 47]), such a simplified setting has limited application value. Consider, for example, those with a visual impairment [16], they will likely struggle to take clean and in-focus images; or can they point to scene-text regions. More recent works [61, 45] begin to study text-based VideoQA which allows QA on multiple images to reduce the proportion of unanswerable cases. Nonetheless, they still assume people know well the locations of scene text (see RoadTextVQA [45] in Fig. 1). Corresponding VQA models, while achieving good results on such benchmarks, cannot reason over user intentions other than an extraction of the OCR results [19].
@@ 2.1 Scene-Text VQA Benchmark
In the scene-text VQA field, various image- and video-based datasets [42, 49, 3, 61, 45, 34, 28] have been proposed. The datasets like TextVQA [42], ST-VQA [3], and ESTVQA [49] provide high-quality images and questions that clearly point to scene text, but such simple settings limit practical applications. To expand the application scope, the datasets like M4viteVQA [61] and RoadTextVQA [45] offer text-rich videos, but design simple questions referring to well-focused scene text. Such simplified settings make QA less practical and they challenge no more than an identification of the scene text. In addition, RoadTextVQA focuses merely on road driving. In this work, we introduce EgoTextVQA to advance research in real-life QA assistance from an ego-perspective. Our videos cover diverse daily scenarios, with the questions reflecting real-user needs yet without the visual attention on scene text.
@@ 2.2 MLLMs for Scene-Text VQA
Recent advancements in Multimodal Large Language Models (MLLMs) [60, 48, 57] have shown significant potential in addressing scene-text VQA tasks [42, 3, 28]. Current MLLMs [29, 60, 23] achieve substantial performance gains in image scene-text QA by enhancing their ability to process higher-resolution images, leading to significant improvements and strong performance. However, these models are limited to image-level input, leaving their video comprehension capabilities largely unexplored. To address this limitation, recent models [53, 58, 47, 25, 8, 18, 7] are developed to support video-level or multi-frame inputs while retain their performance on image-based scene-text QA datasets. Nonetheless, these models have been evaluated separately on image-based OCR-rich benchmarks [42, 3, 28, 34, 32, 33] and video understanding benchmarks [31, 51, 14, 24], lacking suitable assessment for scene-text comprehension in more realistic and dynamic environments. In this work, we will comprehensively analyze the model behaviors in egocentric QA assistance in real-life dynamic environments.
@@ 3 EgoTextVQA Dataset
Figure 2: Examples of EgoTextVQA. Scene text plays pivotal role in understanding and answering the questions which reflect real-user needs. Yet, the videos are without the visual focus on scene text.
@@ 3.1 Dataset Creation
We leverage existing text-rich video sequences and the multimodal understanding and generation capabilities of MLLMs to create our dataset. The generated QA candidates are carefully checked and corrected by human annotators as necessary to ensure the dataset’s high diversity and quality.
@@
Shopping 258/78 Q: Where can I find the “Milano” cookies? A: Top shelf on the right.
@@
Raw Video Filtering Our videos are drawn from two public ego-view video datasets, RoadTextVQA [45] and EgoSchema [31]. They cover outdoor and indoor scenarios, centered around driving and housekeeping activities. RoadTextVQA has 3,222 ten-second videos from dashcam footage of people driving. EgoSchema features 5,063 three-minute videos sourced from Ego4D [15]; its videos are mostly about indoor activities such as cooking, game playing, manufacturing, etc. To find videos rich in scene text, we first apply a state-of-the-art scene-text detection system [17] to the raw videos and threshold for videos with a significant percentage of frames with scene text: exceeding 5% for Ego4D and 15% for RoadTextVQA (since it features more scene text). After automatic filtering, we obtain 700 videos from RoadTextVQA [45] and 1,800 from EgoSchema [15]. Since Ego4D videos are not collected for scene text study, we further filter the unqualified videos, mainly removing the tedious videos and videos that contain watermark or text present only on the clothes of the camera wearers. Finally, we obtain 933/700 videos from Ego4D/RoadTextVQA.
QA Generation Manually collecting QA pairs is labor-intensive; the QAs may also lack diversity. We exploit advanced MLLMs (i.e., GPT-4o) and generate the initial QAs with the following protocol: First, a video is decoded into 6 frame-per-second (fps) and evenly divided into 5 segments, wherein the frames without scene text are removed. Then, we uniformly sample 3 frames from each segment, and feed them to GPT-4o along with the prompts (in Appendix D) to generate 3 QA pairs. This leads to 15 or fewer QA pairs for each video. The prompts are carefully designed to elicit questions that are: (1) goal-oriented, capturing real user needs by engaging with the visual scene from a first-person perspective; (2) aware of scene text, though not necessarily requiring the exact text transcription; (3) naturally expressed in a colloquial, first-person manner with referring expressions; and (4) challenging, as they demand an understanding of the video beyond a single image.
@@
Specifically, we invite 9 well-trained university students for annotation. The manual check has 5 stages: (1) five annotators review for problematic QAs which may be redundant, not reflect real-user needs, and irrelevant to scene text. Where possible, the questions are amended; otherwise, they are removed, leaving half of the original QAs; (2) and (3) two additional annotators re-examine and correct any mechanical issues, further reducing the QAs to 30%; (4) five annotators refine the data to improve its colloquial quality; (5) four annotators strictly refine the remaining QAs according to the criteria specified in QA generation. The participants at each stage are asked to first check the remaining QAs from the prior stage. Additionally, we manually enrich $\sim$10time-sensitive questions to simulate real-time live QA by copying each suitable question to a different timestamp and invite five additional annotators to provide concise answers based on the visual content at the new timestamp. Ultimately, we obtain 7,064 QA pairs of 1,507 videos to form the EgoTextVQA dataset. Detailed question type and video scenario labels are additionally provided by the authors. Figure 2 shows some representative examples; more are presented in Appendix B.5.
@@ 3.2 Dataset Analysis
Dataset Comparison We compare EgoTextVQA with several related benchmarks in Table 2. EgoTextVQA has several unique features. Firstly, it is the first VideoQA testbed designed for egocentric scene-text aware QA assistance in the wild, containing 7K QA pairs that reflect diverse user intentions across 1.5K different egocentric visual scenarios covers both indoor and outdoor activities. The QA pairs focus on scene text comprehension, with approximately 45% of answers and 52% of questions directly referencing the exact scene text. Notably, EgoTextVQA is also the first VideoQA dataset to introduce real-time QA, providing detailed timestamps for each question. While EgoMemoria [54] also includes question timestamps, its answers are based on the entire video content. In contrast, the answers in EgoTextVQA are derived from the video content before the question timestamp and vary according to the specific timestamp. Moreover, the questions in EgoTextVQA emphasize user intention inferring rather than pure visual understanding, and question categories are provided to facilitate model analysis.
@@ 4.1 Experimental Setups
InternVL2 [8] InternLM2.5-7B $448^{2}$ 32 15.8 / 1.4 21.9 / 1.7 14.8 / 1.0 14.5 / 1.2 13.6 / 1.3 16.4 / 1.3
InternVL2 [8] InternLM2-20B $448^{2}$ 16 21.5 / 1.7 27.5 / 2.0 23.5 / 1.5 22.9 / 1.7 22.5 / 1.7 23.5 / 1.7
@@
Evaluation Metric We use GPT-4o mini [36] to evaluate the semantic similarity between the generated and the ground-truth (GT) answers, thus aligning closely with human scoring. Following prior work [30], we assess the generated answers with two metrics: Accuracy (0-100%, the percentages of “yes” answers from evaluator) and Score (0-5, with 5 being the highest match). Concretely, we format the question, predicted answer, and GT answer along with our customized evaluation prompts (in Appendix D) for GPT-4o mini to determine the “yes/no” and score.
Model Evaluation We collect three closed-source API-based models (GPT-4o [35], Gemini 1.5 Flash [39], Gemini 1.5 Pro [39]) and seven advanced open-source MLLMs (MiniCPM-V 2.6 [53], ShareGPT4Video [7], InternVL2 [8], VILA1.5 [25], LLaVA-NeXT-Video [58], CogVLM2-Video [18], Qwen2-VL [47]) for evaluation. The models have reported the state-of-the-art (SOTA) on existing scene-text VQA benchmarks [3, 42, 16], e.g., Qwen2-VL achieves QA accuracy $\sim$85on ST-VQA [3]. More details about these MLLMs and their specific prompts are presented in Appendix B.1 and D.
@@ 4.2 Result Analysis
Tables 3 and  4 show that all models, especially the open-source ones, struggle to perform well on EgoTextVQA. The best results (Gemini Pro 1.5 [39]) are around 33-34% on both EgoTextVQA-Outdoor and EgoTextVQA-Indoor. This exceeds the best-performing open-source models by $\sim$5and $\sim$9on EgoTextVQA-Outdoor and EgoTextVQA-Indoor respectively. The performances vary significantly among the open-source models, with Qwen2-VL [47] and LLaVA-NeXT-Video [58] being the best while MiniCPM-V 2.6 [53] and ShareGPT4Video [7] the worst.
@@
InternVL2 [8] InternLM2.5-7B $448^{2}$ 32 13.5 / 1.2 12.5 / 1.1 14.5 / 1.3 13.3 / 1.2 14.9 / 1.3 12.4 / 1.1 13.6 / 1.2
InternVL2 [8] InternLM2-20B $448^{2}$ 16 20.3 / 1.6 22.4 / 1.8 18.2 / 1.6 25.6 / 1.8 19.2 / 1.6 21.9 / 1.7 21.0 / 1.7
@@
EgoTextVQA-Outdoor Table 3 shows that the MLLMs consistently perform poorly across different question categories. The open-source models particularly fall short in answering “Location” and “Intention Reasoning” types of questions, while the closed-source models GPT-4o and Gemini tend to fail on “Direction” questions. “Location” questions require a nuanced understanding of the spatial positions of targets within the visual scenes, while “Intention Reasoning” questions demand strong inference capabilities of user intentions conditioned on the egocentric visual situation. The poor performance on these questions reveals the severe limitations of existing open-source models in fine-grained visual understanding and reasoning in egocentric scenarios. For the poor behavior of closed-source models on “Direction” questions, a careful observation of the failure predictions shows that models often answer like “unknown scene text” due to their inability to locate or recognize specific scene text (see Figure 4 Top). Overall, Gemini 1.5 Pro [39] wins across most categories, reaching an overall accuracy of 33.4% and a score of 2.0, establishing it as the SOTA model on EgoTextVQA-Outdoor. However, its performance is still 10% behind humans. This results show that even for powerful MLLMs, perceiving scene texts in complex, dynamic, ego-perspective environments is still a challenging and open research topic.
In Figure 5, we additionally analyze the model behaviors on the real-time QA subset of EgoTextVQA-Outdoor. The results show that all models perform much worse on this subset, with the highest accuracy being 20.2% (vs. 33.4% on the full set), highlighting the significant challenge of egocentric live QA. Interestingly, MiniCPM-V 2.6 [53] stands among the Top-2 open-source models for real-time QA despite it being the worst on the full set. Nevertheless, a careful observation of predictions shows that MiniCPM primarily excels at answering unanswerable questions with “I did not find it”, compared with other models which tend to hallucinate the answers. This reflects the models are essentially weak in providing meaningful answers to satisfy user needs.
@@
EgoTextVQA-Indoor Table 4 shows that MLLMs have difficulty in answering user questions in various indoor scenarios. Concretely, the open-source models are especially weak in answering questions in “Kitchen” and “Shopping” activities. Our qualitative analysis (see Figure 4 Bottom) shows that people often ask about the location of ingredients and utensils in the kitchen, where the models tend to provide vague positional information (e.g., “on the left” rather than “on the top shelf on the left”). In shopping, it is particularly challenging for the models to locate a user-specified item among large and densely packed products. Interestingly, we find that ShareGPT4Video [7] performs the worst across all scenarios even though it includes Ego4D videos for training, indicating that it may not be a scene-text aware MLLM. Again, the closed-source models perform relatively better across different visual scenarios.
Intriguingly, we find that humans perform even worse than the closed-source models. Feedback from our human annotators and a careful analysis of the answers suggest that the primary reasons are as follows: (1) scene text recognition is challenging for humans in Ego4D videos due to lower resolution, motion blur, occlusion, and long video; (2) answering the questions needs external knowledge beyond simple scene text recognition. EgoTextVQA-Outdoor mainly involves knowledge about store or restaurant chains (e.g., “Home Deposells building materials), and EgoTextVQA-Indoor focuses on items like kitchen utensils, board games, and specialized books (e.g., identifying amandolineslices potatoes or aquantumfield theory book); (3) answer diversity also results in relatively lower human performance. The first two points underscore the challenges and significance of research on egocentric scene-text QA assistance. The third point unveils a limitation of our dataset, which we wish to solve in the feature by enriching the ground-truth answers for each question. We provide more human study analysis in Appendix B.4 and C.

### v2

@@ 1 Introduction
The advances in wearable cameras and egocentric vision research [12, 14, 28, 39] have led to a surge of interest in developing assistance applications. A promising direction is egocentric video question answering (VideoQA) [55, 26, 58, 15, 11, 62], where AI agents offer live help by answering user questions. Despite the potential, related dataset research [60, 12, 20, 33, 9] has been designed to assess VideoQA models in their visual understanding capabilities. The questions are not curated to reflect real user needs in assistive applications (see EgoTaskQA in Figure 1). QAEgo4D [2] and AssistQ [53] are closer in aim, but these datasets are designed to assist with episodic memory and “how to” demonstration; they emphasize pure visual understanding while neglecting scene text.
Research in assistive techniques [16, 68, 45, 25] have shown that scene text, including signs, labels, and other text elements that are ubiquitous in our daily lives, is frequently involved in user questions when seeking assistance. Existing efforts [45, 69, 3] for QA on scene-text primarily aim at Optical Character Recognition (OCR) from ideal imaging conditions. They assume that people take good pictures and ask questions directly pinpointing the scene-text regions (see TextVQA [45] in Figure 1). Despite the progress (e.g., leading results have accuracies of 85% on ST-VQA [3, 50]), such a simplified setting has limited application value. Consider, for example, those with a visual impairment [16], they will likely struggle to take clean and in-focus images; or can they point to scene-text regions. More recent works [67, 48] begin to study text-based VideoQA which allows QA on multiple images to reduce the proportion of unanswerable cases. Nonetheless, they still assume people know well the locations of scene text (see RoadTextVQA [48] in Figure 1). Corresponding VQA models, while achieving good results on such benchmarks, cannot reason over user intentions other than an extraction of the OCR results [19].
@@ 2.1 Scene-Text VQA Benchmark
In the scene-text VQA field, various image- and video-based datasets [45, 52, 3, 67, 48, 36, 30] have been proposed. The datasets like TextVQA [45], ST-VQA [3], and ESTVQA [52] provide high-quality images and questions that clearly point to scene text, but such simple settings limit practical applications. To expand the application scope, the datasets like M4viteVQA [67] and RoadTextVQA [48] offer text-rich videos, but they design simple questions which refer to well-focused scene text. Such simplified settings make QA less practical and they challenge no more than an identification of the scene text. In addition, RoadTextVQA focuses merely on road driving. In this work, we introduce EgoTextVQA to advance research in real-life QA assistance from an egocentric perspective. Our videos cover diverse daily scenarios, with the questions reflecting real user needs yet without the visual attention on scene text.
@@ 2.2 MLLMs for Scene-Text VQA
Recent advancements in Multimodal Large Language Models (MLLMs) [66, 51, 63] have shown significant potential in addressing scene-text VQA tasks [45, 3, 30]. Current MLLMs [31, 66, 23] achieve substantial performance gains in image scene-text QA by enhancing their ability to process higher-resolution images, leading to significant improvements and strong performance. However, these models are limited to image-level input, leaving their video comprehension capabilities largely unexplored. To address this limitation, recent models [59, 64, 50, 27, 7, 57, 56, 18, 6] are developed to support video-level or multi-frame inputs while retain their performance on image-based scene-text QA datasets. Nonetheless, these models have been evaluated separately on image-based OCR-rich benchmarks [45, 3, 30, 36, 34, 35] and video understanding benchmarks [33, 54, 13, 24, 43], lacking suitable assessment for scene-text comprehension in more realistic and dynamic environments. In this work, we will comprehensively analyze the model behaviors in egocentric QA assistance in real-life dynamic environments.
@@ 3 EgoTextVQA Dataset
Figure 2: Examples of EgoTextVQA. Scene text plays pivotal role in understanding and answering the questions which reflect real user needs. Yet, the videos are without the visual focus on scene text.
@@ 3.1 Dataset Creation
We leverage existing text-rich video sequences and the multimodal understanding and generation capabilities of MLLMs to create our dataset. The generated QA candidates are carefully reviewed and refined by human annotators as necessary to ensure the dataset’s high diversity and quality.
@@
Shopping 358/78 Q: Where can I find the “Milano” cookies? A: Top shelf on the right.
@@
Raw Video Filtering Our videos are drawn from two public ego-view video datasets, RoadTextVQA [48] and EgoSchema [33]. They cover outdoor and indoor scenarios, centered around driving and housekeeping activities. RoadTextVQA has 3,222 ten-second videos from dashcam footage of people driving. EgoSchema features 5,063 three-minute videos sourced from Ego4D [14]; its videos are mostly about indoor activities such as cooking, game playing, manufacturing, etc. To find videos rich in scene text, we first apply a state-of-the-art scene-text detection system [17] to the raw videos and threshold for videos with a significant percentage of frames with scene text: exceeding 5% for Ego4D and 15% for RoadTextVQA (since it features more scene text). After automatic filtering, we obtain 700 videos from RoadTextVQA [48] and 1,800 from EgoSchema [14]. Since Ego4D videos are not collected for scene text study, we further filter the unqualified videos, mainly removing the tedious videos and videos that contain watermark or text present only on the clothes of the camera wearers. Here, we obtain 933/700 videos from Ego4D/RoadTextVQA.
QA Generation Manually collecting QA pairs is labor-intensive; the QAs may also lack diversity. We use advanced MLLMs (i.e., GPT-4o) and generate the initial QAs with the following protocol: First, a video is decoded into 6 frames per second (fps) and evenly divided into 5 segments, wherein the frames without scene text are removed. Then, we uniformly sample 3 frames from each segment, and feed them to GPT-4o along with the prompts (in Appendix D) to generate 3 QA pairs. This leads to a maximum of 15 QA pairs per video. The prompts are carefully designed to elicit questions that are: (1) goal-oriented, capturing real user needs by engaging with the visual scene from a first-person perspective; (2) aware of scene text, though not necessarily requiring the exact text transcription; (3) naturally expressed in a colloquial, first-person manner with referring expressions; and (4) challenging, as they demand an understanding of the video beyond a single image.
@@
Specifically, we invite 9 well-trained university students for annotation. The manual check has 5 stages: (1) five annotators review for problematic QAs which may be redundant, not reflect real user needs, and irrelevant to scene text. The questions are amended where possible, or removed, leaving half of the original QAs; (2) and (3) two additional annotators re-examine and correct any mechanical issues, further reducing the QAs to 30%; (4) five annotators refine the data to improve its colloquial quality; (5) four annotators strictly refine the remaining QAs according to the criteria specified in QA generation. The participants at each stage are asked to first check the remaining QAs from the prior stage. Additionally, we manually enrich $\sim$10time-sensitive questions to simulate real-time live QA by copying each suitable question to a different timestamp and invite five additional annotators to provide concise answers based on the visual content at the new timestamp. Ultimately, we obtain 7,064 QA pairs of 1,507 videos to form the EgoTextVQA dataset. Detailed question type and video scenario labels are provided. Figure 2 shows some representative examples; more are presented in Appendix B.5.
@@ 3.2 Dataset Analysis
Dataset Comparison We compare EgoTextVQA with several related benchmarks in Table 2. EgoTextVQA has several unique features. Specifically, EgoTextVQA is the first VideoQA testbed designed for egocentric scene-text aware QA assistance in real-world scenarios, containing 7K QA pairs across 1.5K egocentric visual scenarios, covering both indoor and outdoor activities. The QA pairs focus on scene text comprehension, with about 45% of answers and 52% of questions referencing the exact scene text. Notably, EgoTextVQA introduces real-time QA, providing detailed timestamps for each question. While EgoMemoria [60] also includes timestamps, its answers are based on the entire video, whereas EgoTextVQA answers are derived from video content before the question timestamp. Additionally, the questions in EgoTextVQA focus on inferring user intentions rather than pure visual understanding, with question categories provided for model analysis.
@@ 4.1 Experimental Setups
InternVL2-8B [7] InternLM2.5-7B $448^{2}$ 32 15.8 / 1.4 21.9 / 1.7 14.8 / 1.0 14.5 / 1.2 13.6 / 1.3 16.4 / 1.3
InternVL2-26B [7] InternLM2-20B $448^{2}$ 16 21.5 / 1.7 27.5 / 2.0 23.5 / 1.5 22.9 / 1.7 22.5 / 1.7 23.5 / 1.7
@@
Evaluation Metric We use GPT-4o mini [38] to evaluate the semantic similarity between the generated and the ground-truth (GT) answers, thus aligning closely with human scoring. Following [32], we assess the generated answers with two metrics: Accuracy (0-100%, the percentage of “yes” answers from the evaluator) and Score (0-5, with 5 being the highest match). Concretely, we format the question, predicted answer, and GT answer along with our customized evaluation prompts (in Appendix D) for GPT-4o mini to determine the “yes/no” and score.
Model Evaluation We evaluate three closed-source API-based models (GPT-4o [37], Gemini 1.5 Flash [41], Gemini 1.5 Pro [41]) and seven advanced open-source MLLMs (MiniCPM-V 2.6 [59], ShareGPT4Video [6], InternVL2 [7], VILA1.5 [27], LLaVA-NeXT-Video [64], CogVLM2-Video [18], Qwen2-VL [50]) for evaluation. The models have reported the state-of-the-art (SOTA) on existing scene-text VQA benchmarks [3, 45, 16], e.g., Qwen2-VL achieves QA accuracy $\sim$85on ST-VQA [3]. More details about these MLLMs and their specific prompts are presented in Appendix B.1 and  D.
@@ 4.2 Result Analysis
Tables 3 and 4 show that all models, especially the open-source ones, struggle to perform well on EgoTextVQA. The best results, achieved by Gemini Pro 1.5 [41], are around 33$\sim$34on both EgoTextVQA-Outdoor and EgoTextVQA-Indoor, exceeding the best-performing open-source models by approximately $\sim$5and $\sim$9on EgoTextVQA-Outdoor and EgoTextVQA-Indoor, respectively. The performances vary significantly among open-source models, with Qwen2-VL [50] and LLaVA-NeXT-Video [64] performing the best, while MiniCPM-V 2.6 [59] and ShareGPT4Video [6] perform the worst.
@@
InternVL2-8B [7] InternLM2.5-7B $448^{2}$ 32 13.5 / 1.2 12.5 / 1.1 14.5 / 1.3 13.3 / 1.2 14.9 / 1.3 12.4 / 1.1 13.6 / 1.2
InternVL2-26B [7] InternLM2-20B $448^{2}$ 16 20.3 / 1.6 22.4 / 1.8 18.2 / 1.6 25.6 / 1.8 19.2 / 1.6 21.9 / 1.7 21.0 / 1.7
@@
EgoTextVQA-Outdoor Table 3 shows that MLLMs struggle across different question categories. The open-source models perform particularly poorly on “Location” and “Intention Reasoning” questions, while closed-source models, GPT-4o and Gemini 1.5 Pro/Flash, often fail on “Direction” questions. “Location” questions require understanding spatial relationships, while “Intention Reasoning” involves inferring user intent from egocentric visual context. These challenges highlight the limitations of open-source models in fine-grained visual reasoning. For “Direction” questions, closed-source models frequently respond with phrases like “unknown scene text” suggesting difficulties in recognizing or locating specific scene text (see Figure 4 Top). Overall, Gemini 1.5 Pro [41] performs best, achieving 33.4% accuracy and a score of 2.0, making it the SOTA model on EgoTextVQA-Outdoor. However, it still trails human performance by $\sim$10%, indicating that even advanced MLLMs struggle with scene text perception in complex, dynamic, egocentric environments.
In Figure 5, we further analyze model behavior on the real-time QA subset of EgoTextVQA-Outdoor. The results show that all models perform significantly worse on this subset, with the highest accuracy reaching only 20.2% (vs. 33.4% on the full set), underscoring the substantial challenge of egocentric live QA. Interestingly, MiniCPM-V 2.6 [59] ranks among the top two open-source models for real-time QA, despite being the worst on the full set. However, a closer examination of its predictions reveals that MiniCPM-V 2.6 primarily excels at answering unanswerable questions with responses like “I did not find it” whereas other models tend to hallucinate answers. This finding suggests that current models remain fundamentally weak in generating meaningful answers that satisfy user needs.
@@
EgoTextVQA-Indoor Table 4 shows that MLLMs struggle with the questions across indoor scenarios, especially in “Kitchen” and “Shopping” activities. Our qualitative analysis (see Figure 4 Bottom) shows that people often ask about the location of ingredients and utensils in the kitchen, where the models tend to provide vague positional information (e.g., “on the left” rather than “on the top shelf on the left”). In shopping, it is challenging for the models to locate a user-specified item among large and densely packed products. ShareGPT4Video [6] performs the worst despite training on Ego4D videos, suggesting limited scene-text awareness. Meanwhile, closed-source models perform relatively better across scenarios.
Intriguingly, we find that humans perform even worse than the closed-source models. Feedback from our human annotators and a careful analysis of the answers suggest that the primary reasons are as follows: (1) scene text recognition is challenging for humans in Ego4D videos due to lower resolution, motion blur, occlusion, and long video; (2) answering the questions needs external knowledge beyond simple scene text recognition. EgoTextVQA-Outdoor mainly involves knowledge about store or restaurant chains (e.g., “Home Depo” sells building materials), and EgoTextVQA-Indoor focuses on items like kitchen utensils, board games, and specialized books (e.g., identifying a “mandoline” slices potatoes or a “quantum” field theory book); (3) answer diversity also results in relatively lower human performance. The first two points underscore the challenges and significance of research on egocentric scene-text QA assistance. The third point unveils a limitation of our dataset, which we wish to solve in the future by enriching the GT answers for each question. We provide more human study analysis in Appendix B.4 and  C.